1/2

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,
Varulven  The Werewolf  La lupfantomo  Der Werwolf

translated by Sven Collberg translated by Alexander Gross translated by Bertilo Wennergren

En varulv lopp från fru m.m. en vacker afton samt begav sig till en bymagisters grav och bad: “Var snäll och konjugera.

A Werewolf, troubled by his name, Left wife and brood one night and came To a hidden graveyard to enlist The aid of a long-dead philologist.

Jen lupfantom’ de famili’ meznokte iam kuris for al tombo de la profesor’. kaj petis pri konjugaci’.

Ein Werwolf eines Nachts entwich von Weib und Kind und sich begab an eines Dorfschullehrers Grab und bat ihn: �Bitte, beuge mich!“

mig!” Byskolläraren stod upp på blecknamnsskyltens

mässingknopp med svar till den, som satt på pass med tåligt korslagd tass på tass.

“Oh sage, wake up, please don’t berate me,”

He howled sadly, “Just conjugate me.”

The seer arose a bit unsteady

Yawned twice, wheezed once, and then was ready.

Do diktis tiu dokta ul’ starante sur la kructabul’ dum lup’ senmova kiel splinto atentis vortojn de l’ mortinto.

Der Dorfschulmeister stieg hinauf auf seines Blechschilds Messing-

knauf und sprach zum Wolf, der seine Pfo-

ten geduldig kreuzte vor dem Toten:

“Jag var-ulv 1:a pers. i sing., du, han, hon, den, det var-ulv. Så vi voro-ulv, I voren. På de voro-ulv kom ingenting.”

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

De “lupfantomo”, simpla rekta, tra “lupfintomo”, plej perfekta, plu “lupfontomo”, iom rara, ĝis “lupfuntomo”, ho, koŝmara.

�Der Werwolf“ – sprach der gute

Mann, �des Werwolfs, Genitiv sodann, dem Werwolf, Dativ, wie man’s

nennt, den Werwolf, - damit hat’s ein End!“

Varulven myste gladåt hur pers. hade följts av sing. och plur. “Fast”, bad han, “kan det inte ges

ens en enda pers. i tempus presens?”

“I know that–I’m no mental cripple–

The future form and participle

Are what I crave,” the beast replied. The scholar paused–again he tried:

Ĉi participoj plaĉis multe al la fantomo, kaj rezulte ĝi petis: “Legu plu el PIV-o pri participoj en pasivo.”

Dem Werwolf schmeichelten die

Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl

noch!“

…  …  …  …

2/2

Bekänna måste vår magister, att därvidlag fanns vissa brister. En varulv sine imperfecto var contradictio in adjecto.

“A ‘Will-be-wolf?’ It’s just too long:

‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly

wrong! Such words are wounds beyond all suture– I’m sorry, but you have no future.”

La viro ĉagreniĝis tamen:

“Jam fermis mi per fin’ kaj

amen, ĉar netransiraj bestoj feblas kaj la pasiv’ neniam eblas.”

Der Dorfschullehrer aber musste gestehn, dass er von ihr nichts wus-

ste. Zwar Wölfe gab’s in großer Schar, doch �Wer“ gab’s nur im Sinular.

Varulven satt med tårad lins. Han visste att han fanns och finns. Därtill kom också fru m.m. som även måste existera.

The Werewolf knew better–his sons still slept At home, and homewards now he crept, Happy, humble, without apology For such folly of philology.

Sed lupfantomo pri transiro ja spertas pli ol eĉ vampiro. La lup’ revenis larmokula al sia famili’ ulula.

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und Kind! Doch da er keine Gelehrter eben, so schied er dankend und ergeben.

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into Swedish by Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Translation of the German poem “Der Werwolf”

by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow.

com/ literaturo/ lupfantomo. html .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into Esperanto by Bertilo Wennergren.

Arg-825-1653 (2013-02-14 23:29:47)

Vidu la retejon http: // bertilow. com/ literaturo/ lupfantomo. html .

Author of this German poem is Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)